# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ Abstract
This paper studies the gradient flow dynamics that arise when training deep homogeneous neural networks, starting with small initializations. The present work considers neural networks that are assumed to have locally Lipschitz gradients and an order of homogeneity strictly greater than two. This paper demonstrates that for sufficiently small initializations, during the early stages of training, the weights of the neural network remain small in norm and approximately converge in direction along the Karush-Kuhn-Tucker (KKT) points of the neural correlation function introduced in [1]. Additionally, for square loss and under a separability assumption on the weights of neural networks, a similar directional convergence of gradient flow dynamics is shown near certain saddle points of the loss function.
@@ 1 Introduction
Neural networks have achieved remarkable success across various tasks, yet the precise mechanism driving this success remains theoretically elusive. The training of neural networks involve optimizing a non-convex loss function, where the training algorithm typically is a first-order methods such as gradient descent or its variants. A particularly puzzling aspect is how these training algorithms succeed in finding a solution with good generalization capabilities despite the non-convexity of the loss landscape. In addition to the choice of the training algorithm, the choice of initialization in these algorithms play a crucial role in determining the neural network performance. Indeed, recent works have made increasingly clear the benefit of small initializations, revealing that neural networks trained using (stochastic) gradient descent with small initializations exhibit feature learning [2] and also generalize better for various tasks [3, 4, 5]; see Section 2 for more details into the impact of initialization scale. However, for small initializations, the training dynamics of neural networks is extremely non-linear and not well understood so far. Our focus in this paper is on understanding the effect of small initialization on the training dynamics of neural networks. In pursuit of a deeper understanding of the training mechanism for small initializations, researchers have uncovered the phenomenon of directional convergence in the neural network weights during early phases of training [6, 7]. In [6], authors study the gradient flow dynamics of training two-layer Rectified Linear Unit (ReLU) neural networks. They demonstrate that in the early stages of training, the weights of two-layer ReLU neural networks converge in direction while their norms remain small. This phenomenon is referred to as early directional convergence. In a recent work [1], early directional convergence was further established for two-homogeneous [^1] A neural network $\mathcal{H}$ is defined to be $L$-(positively) homogeneous, if its output $\mathcal{H}(\mathbf{x};\mathbf{w})$ satisfies $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),\forall c\geq 0,$ where $\mathbf{x}$ is the input, $\mathbf{w}$ is a vector containing all the weights, and $L$ is referred to as the order of homogeneity. neural networks, illustrating the importance of homogeneity for such types of phenomenon (as opposed to the specific choice of activation function or architecture). Despite focusing solely on the early phases of training, these works demonstrate that interesting behaviors begin to emerge very early in the training dynamics. Moreover, these insights in the initial stages of training have been useful towards a comprehensive understanding of the entire training dynamics in some special cases [8, 9, 10]. So far, all the works on the early stages of training dynamics of neural networks for small initialization have only considered two-layer neural networks (or more broadly two-homogeneous). In contrast, our understanding for deep neural networks is much more limited. In this work, we take a first step towards this goal by focusing on deep homogeneous neural networks (with locally Lipschitz gradients) and establish early directional convergence for them. To our knowledge, there has been no prior work that rigorously demonstrates early directional converge for deeper neural networks. To describe the phenomenon of early directional convergence, [1] introduced the notion of Neural Correlation Function (NCF). For a given neural network and a vector, the NCF quantifies the correlation between the output of neural network and the vector; we provide a more precise definition in Section 4. They show that in the early stages of training, the weights of the neural network remain small and converge in direction to a KKT point of the constrained NCF, an optimization problem that maximizes the NCF on unit sphere. Furthermore, they also establish similar directional convergence near certain saddle points as well. We emphasize that their results are specific to two-homogeneous neural networks. The present work studies the phenomenon of early directional convergence for deep homogeneous neural networks. We particularly consider $L$-homogeneous neural networks, where $L>2$, that have locally Lipschitz gradients. Our contribution can be summarized as follows:
@@
  • In Theorem 5.1 we describe the gradient flow dynamics of deep homogeneous neural networks, with order of homogeneity strictly greater than two, trained with small initializations. Specifically, for square and logistic losses, we show that if the initialization is sufficiently small, then in the early stages of training the weights remain small and either approximately converge in direction to a non-negative KKT point of the constrained NCF defined with respect to the labels in the training set, or are approximately $\mathbf{0}$.
  • For square loss, Theorem 5.3 establishes a similar directional convergence for deep homogeneous neural networks near certain saddle points, provided the weights of neural network have a separable structure. We show that when initialized in a sufficiently small neighborhood of a saddle point, then under gradient flow the weights remain near the saddle point for a sufficiently long time, during which the weights with small norm remain small and either approximately converge in direction along non-negative KKT points of the constrained NCF defined with respect to the residual error at that saddle point, or are approximately $\mathbf{0}$.
@@
To prove our results, we build upon findings from [1] along with crucial new insights; see Section 5.1.1 for details. It is worth noting that we assume neural networks to have locally Lipschitz gradients, which excludes deep ReLU networks. However, it does include deep linear networks [11, 12], and deep neural networks with differentiable and homogeneous activation such as higher powers of ReLU $\phi(x):=\max(x,0)^{p}$ for $p\geq 2$ [13, 14, 15] and polynomials [16, 17]. Moreover, we also discuss the challenges in extending our results for deep ReLU networks, which could be useful in extending our results to ReLU networks.
@@ 2 Related Works
Understanding the training dynamics of neural networks have been the subject of numerous investigations [18, 4, 19, 20, 21]. These investigations have revealed two contrasting viewpoints based on the scale of initialization. For large initialization regime, the training dynamics of wide neural networks are effectively captured by a kernel referred to as Neural Tangent Kernel (NTK) [18]. In this regime, the weights remain close to initialization during training, and neural networks behaves like the linearization of neural network around its initialization [4]. On the other hand, in small initialization regime, training dynamics is extremely non-linear, and the change in weights during training is significant [21, 22, 3, 2]. Moreover, studies like [3, 4] have observed improved generalization with decreasing initialization scale for various tasks, which makes understanding the dynamics of neural networks in the small initialization regime crucial for better understanding of practical success of deep networks. While considerable progress has been made on understanding the dynamics of neural networks in the large initialization regime, the theoretical investigations into the small initialization regime has been comparatively limited, primarily due to the extremely non-linear dynamics during training. Such investigations have predominantly focused on linear neural networks [5, 12, 23] and shallow non-linear neural networks [22, 21, 24, 20]. Even for these networks, a comprehensive understanding of the gradient descent dynamics is further limited to diagonal linear networks [5], shallow matrix factorization [25, 26], and two-layer (Leaky) ReLU networks under various simplifying assumptions on the datasets [8, 9, 27, 10, 28]. Our work could be the first step towards a greater understanding of the training dynamics of deep neural networks. The phenomenon of early directional convergence for small initializations was first investigated by [6] for two-layer ReLU neural networks. This was later extended to two-homogeneous neural networks by [1], which includes, in addition to two-layer ReLU neural networks, deep ReLU neural networks with only two trainable layers and single layer squared ReLU neural networks. Other empirical investigations have observed early directional convergence in some deeper networks as well. For example, [29] looks at the early training dynamics of three-layer ReLU neural networks and observe that weights converge along some specific orientations. In [30] authors empirically show a alignment effect in the kernel of the neural network during early stages of the training. However, rigorous investigations thus far have been limited to shallow neural networks. For classification tasks, another line of work studies directional convergence of weights in the later stages of training homogeneous neural networks [31, 32]. In these works, apart from homogeneity, the tail behavior of the loss functions typically used for classification, for example exponential and logistic loss, also plays an important role. Although the problem considered in this paper is different, some theoretical results developed in these works were helpful for us.
3 Preliminaries
We briefly state important notations and describe some relevant preliminary concepts. For any $N\in\mathbb{N}$, we let $[N]=\{1,2,\ldots,N\}$ denote the set of positive integers less than or equal to $N$. We let $\|\cdot\|_{2}$ denote the $\ell_{2}$ norm for a vector and the spectral norm for a matrix. We write $\dot{x}(t)=:\frac{dx(t)}{dt}$, and for the sake of brevity we may remove the independent variable $t$ if it is clear from context. A $k$-dimensional sphere is denoted by $\mathcal{S}^{k-1}$. Suppose $P$ and $Q$ are locally compact metric spaces, such as Euclidean spaces. Then, a function $f:P\rightarrow Q$ is called locally Lipschitz if and only if it is Lipschitz continuous on every compact subset of $Q$. We will assume the neural networks consider in this paper are definable under some o-minimal structure. This is a mild technical assumption and includes deep neural networks with common activation functions such as ReLU, higher powers of ReLUs and polynomials. Such assumptions have mainly been made in the past works, and here as well, to ensure convergence of certain gradient flow trajectories [31].
@@ 4 Problem formulation
We adopt a supervised learning framework for training, where we assume $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$ is the training dataset, and let $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]\in\mathbb{R}^{d\times n}$ and $\mathbf{y}=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n}$. For a neural network $\mathcal{H}$, its output is denoted by $\mathcal{H}(\mathbf{x};\mathbf{w})$, where $\mathbf{x}$ is the input and $\mathbf{w}\in\mathbb{R}^{k}$ is the vector containing all the weights. We let $\mathcal{H}(\mathbf{X};\mathbf{w})=\left[\mathcal{H}(\mathbf{x}_{1};\mathbf{w}),\ldots,\mathcal{H}(\mathbf{x}_{n};\mathbf{w})\right]^{\top}\in\mathbb{R}^{n}$ be the vector containing the output of neural network for all inputs, and $\mathcal{J}(\mathbf{X};\mathbf{w}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ denotes the Jacobian of $\mathcal{H}(\mathbf{X};\mathbf{w})$ with respect to $\mathbf{w}$. The following assumptions formalizes the previously mentioned properties of neural networks considered in this paper.
@@
We make the following tripartite assumption on the neural network function: (i) For any fixed $\mathbf{x}$, $\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz and is definable under some o-minimal structure that includes polynomials and exponential222 [33] proved the existence of such an o-minimal structure.. (ii) For all $c>0$, $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),$ for some $L>2$ (i.e., the network is $L$-homogenous). (iii) The gradient [^3] Throughout the paper, $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$ denotes the gradient of $\mathcal{H}(\mathbf{x};\mathbf{w})$ with respect to $\mathbf{w}$. $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz.
@@
where $\ell(\hat{y},y)$ is the loss function and it is either the square loss, $\ell(\hat{y},y)=(\hat{y}-y)^{2}/2,$ or the logistic loss, $\ell(\hat{y},y)=2\ln(1+e^{-\hat{y}y}).$ We minimize (1) using gradient flow, which leads to following differential equation
@@
where $\delta$ is a positive scalar that controls the scale of initialization, and $\mathbf{w}_{0}$ is a vector. Following [1], for a fixed vector $\mathbf{z}\in\mathbb{R}^{n}$ and neural network $\mathcal{H}$, the Neural Correlation Function (NCF) is defined here as
@@
and the constrained optimization problem
$\max_{\|\mathbf{w}\|^{2}_{2}=1}\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{w})$
is referred to as the constrained NCF.
5 Main Results
5.1 Early Directional Convergence
Our first main result is the following theorem which establishes the approximate directional convergence in the early stages of training $L$-homogeneous neural networks with small initializations, where $L>2$.
Theorem 5.1.
Let $\mathbf{w}_{0}$ be a fixed unit vector. For any $\epsilon\in(0,\eta/2)$, where $\eta$ is a positive constant, there exists $T_{\epsilon},\tilde{B}_{\epsilon}$ and $\overline{\delta}>0$ such that the following holds: for any $\delta\in(0,\overline{\delta})$ and solution $\mathbf{w}(t)$ of (2) for square or logistic loss with initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, we have
$\left\|\mathbf{w}\left({t}\right)\right\|_{2}\leq\tilde{B}_{\epsilon}\delta,\text{ for all }t\in\left[0,\frac{T_{\epsilon}}{\delta^{L-2}}\right].$

### v2

@@ Abstract
This paper studies the gradient flow dynamics that arise when training deep homogeneous neural networks assumed to have locally Lipschitz gradients and an order of homogeneity strictly greater than two. It is shown here that for sufficiently small initializations, during the early stages of training, the weights of the neural network remain small in (Euclidean) norm and approximately converge in direction to the Karush-Kuhn-Tucker (KKT) points of the recently introduced neural correlation function. Additionally, this paper also studies the KKT points of the neural correlation function for feed-forward networks with (Leaky) ReLU and polynomial (Leaky) ReLU activations, deriving necessary and sufficient conditions for rank-one KKT points.
@@ 1 Introduction
Neural networks have achieved remarkable success across various tasks, yet the precise mechanism driving this success remains theoretically elusive. The training of neural networks involves optimizing a non-convex loss function, where the training algorithm typically is a first-order method such as gradient descent or its variants. A particularly puzzling aspect is how these training algorithms succeed in finding a solution with good generalization capabilities despite the non-convexity of the loss landscape. In addition to the choice of the training algorithm, the choice of initialization in these algorithms plays a crucial role in determining the neural network performance. Indeed, recent works have made increasingly clear the benefit of small initializations, revealing that neural networks trained using (stochastic) gradient descent with small initializations exhibit feature learning (Yang & Hu, 2021) and also generalize better for various tasks (Chizat et al., 2019; Geiger et al., 2020; Woodworth et al., 2020); see Section 2 for more details into the impact of initialization scale. However, for small initializations, the training dynamics of neural networks is extremely non-linear and not well understood so far. Our focus in this paper is on understanding the effect of small initialization on the training dynamics of neural networks.
In pursuit of a deeper understanding of the training mechanism for small initializations, researchers have uncovered the phenomenon of directional convergence in the neural network weights during the early phases of training (Maennel et al., 2018; Luo et al., 2021). The authors of Maennel et al. (2018) study the gradient flow dynamics of training two-layer Rectified Linear Unit (ReLU) neural networks, and demonstrate that in the early stages of training, the weights of two-layer ReLU neural networks converge in direction while their norms remain small. This phenomenon is referred to as early directional convergence. In a recent work, Kumar & Haupt (2024) further established early directional convergence for two-homogeneous [^1] A neural network $\mathcal{H}$ is defined to be $L$-(positively) homogeneous, if its output $\mathcal{H}(\mathbf{x};\mathbf{w})$ satisfies $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),\forall c\geq 0,$ where $\mathbf{x}$ is the input and $\mathbf{w}$ is a vector containing all the weights. neural networks, illustrating the importance of homogeneity for such types of phenomenon (as opposed to the specific choice of activation function or architecture).
To describe the phenomenon of early directional convergence for two-homogeneous neural networks, Kumar & Haupt (2024) introduced the notion of the Neural Correlation Function (NCF). For a given neural network and a vector, the NCF quantifies the correlation between the output of neural network and the vector; we provide a more precise definition in Section 3. They show that in the early stages of training, the weights of the neural network remain small and converge in direction to a KKT point of the constrained NCF, an optimization problem that maximizes the NCF on unit sphere.
So far, all the works on the early stages of training dynamics of neural networks for small initialization have only considered two-layer neural networks (or more broadly, two-homogeneous networks). In contrast, our understanding for deep neural networks is much more limited. In the first part of this work, we establish early directional convergence for deep homogeneous neural networks. We particularly consider $L$-homogeneous neural networks, where $L>2$, that have locally Lipschitz gradients, and our main result can be summarized as follows:
@@
  • In Theorem 1 we describe the gradient flow dynamics of deep homogeneous neural networks, with order of homogeneity strictly greater than two, trained with small initializations. Specifically, for square and logistic losses, we show that if the initialization is sufficiently small, then in the early stages of training the weights remain small and either approximately converge in direction to a non-negative KKT point of the constrained NCF defined with respect to the labels in the training set, or are approximately zero.
@@
It is worth noting that we assume neural networks to have locally Lipschitz gradients, which excludes deep ReLU networks. However, it does include deep linear networks (Arora et al., 2019a; Saxe et al., 2014), and deep neural networks with differentiable and homogeneous activation such as polynomial ReLU $\phi(x):=\max(x,0)^{p}$ for $p\geq 2$ (Gribonval et al., 2022; Klusowski & Barron, 2018; Zhong et al., 2017) and monomials (Livni et al., 2014; Soltanolkotabi et al., 2019). We also discuss the challenges in extending our results for deep ReLU networks in Appendix D.
The second part of our paper studies the KKT points of the constrained NCF for feed-forward homogeneous neural networks. For such networks, we observe empirically that, along with converging towards a KKT point of the constrained NCF during the early stages of training, the hidden weights also exhibit a rank one structure. Motivated by this observation, we mathematically characterize the rank-one KKT points of feed-forward homogeneous neural networks. Our second main result can be summarized as follows:
  • In Section 5, we derive sufficient conditions for a rank-one KKT point of the constrained NCF for feed-forward neural network with activation function $\sigma(x)=\max(x,\alpha x)^{p}$, for some $p\in\mathbb{N}$, $\alpha\in{\mathbb{R}}$ and arbitrary training data. We also provide matching necessary conditions under additional assumptions and validate these assumptions empirically. Our results also provide a method to obtain rank-one KKT points of the constrained NCF using KKT points of a smaller optimization problem.
@@ 2 Related Works
Understanding the training dynamics of neural networks has been the subject of numerous investigations, including (Arora et al., 2019b; Chizat & Bach, 2020; Chizat et al., 2019; Jacot et al., 2018; Mei et al., 2019). These investigations have revealed two contrasting viewpoints depending on the scale of initialization. For large initialization regimes, the training dynamics of wide neural networks are effectively captured by a kernel referred to as Neural Tangent Kernel (NTK) Jacot et al. (2018). In this regime, the weights remain close to initialization during training, and neural networks behave like their linearizations around their initializations (Chizat et al., 2019). On the other hand, in small initialization regimes, the training dynamics are extremely non-linear, and the change in weights during training is significant (Chizat & Bach, 2018; Geiger et al., 2020; Mei et al., 2019; Yang & Hu, 2021). Moreover, studies like Chizat et al. (2019); Geiger et al. (2020) have observed improved generalization with decreasing initialization scale for various tasks, which makes understanding the dynamics of neural networks in the small initialization regime crucial for better understanding of practical success of deep networks.
While considerable progress has been made on understanding the dynamics of neural networks in the large initialization regime, the theoretical investigations into the small initialization regime have been comparatively limited, primarily due to the extremely non-linear dynamics during training. Such investigations have predominantly focused on linear neural networks (Arora et al., 2019a; Gunasekar et al., 2017; Woodworth et al., 2020) and shallow non-linear neural networks (Chizat & Bach, 2018; 2020; Mei et al., 2019; Rotskoff & Vanden-Eijnden, 2018). Even for these networks, a comprehensive understanding of the gradient descent dynamics is further limited to diagonal linear networks (Woodworth et al., 2020), shallow matrix factorization (Jin et al., 2023; Stöger & Soltanolkotabi, 2021), and two-layer (Leaky) ReLU networks under various simplifying assumptions on the datasets (Boursier et al., 2022; Brutzkus & Globerson, 2019; Lyu et al., 2021; Min et al., 2024; Wang & Ma, 2023). Our work could be the first step towards a greater understanding of the training dynamics of deep neural networks.
The phenomenon of early directional convergence for small initializations was first investigated by Maennel et al. (2018) for two-layer ReLU neural networks. This was later extended to two-homogeneous neural networks by Kumar & Haupt (2024), which includes, in addition to two-layer ReLU neural networks, single layer squared ReLU neural networks and deep ReLU neural networks with only two trainable layers. Other empirical investigations have observed early directional convergence in some deeper networks as well. For example, Zhou et al. (2022) looks at the early training dynamics of three-layer ReLU neural networks and observe that weights converge along some specific orientations. Also, Atanasov et al. (2022) empirically show an alignment effect in the kernel of the neural network during early stages of the training. However, rigorous investigations thus far have been limited to shallow neural networks. Despite focusing solely on the early phases of training, these works demonstrate that interesting behaviors begin to emerge very early in the training dynamics. Moreover, these insights into the initial stages of training have been useful towards a comprehensive understanding of the entire training dynamics in some special cases (Boursier et al., 2022; Min et al., 2024; Lyu et al., 2021; Wang & Ma, 2023).
@@ 3 Problem Formulation
Notation: For any $N\in\mathbb{N}$, we let $[N]=\{1,2,\ldots,N\}$ denote the set of positive integers less than or equal to $N$. We let $\|\cdot\|_{2}$ denote the $\ell_{2}$ norm for a vector and the spectral norm for a matrix. For a vector $\mathbf{z}\in\mathbb{R}^{n}$, $z_{i}$ denotes its $i$th entry, $|\mathbf{z}|=[|z_{1}|,|z_{2}|,\cdots,|z_{n}|]^{\top}$, and $\mathbf{z}^{q}=[z_{1}^{q},z_{2}^{q},\cdots,z_{n}^{q}]^{\top}$, where $q\in{\mathbb{N}}$. We write $\dot{x}(t)=:\frac{dx(t)}{dt}$, and for the sake of brevity we may remove the independent variable $t$ if it is clear from context. For any locally Lipschitz continuous function $f:X\to{\mathbb{R}}$, its Clarke sub-differential is denoted by $\partial f({\mathbf{x}})$. For $\sigma:{\mathbb{R}}\rightarrow{\mathbb{R}}$ and $p\in{\mathbb{N}}$, $\sigma^{p}(\cdot)$ denotes the function resulting from composing $\sigma(\cdot)$ with itself $p$ times, and we assume $\sigma^{0}(x)=x$. We define a KKT point of an optimization problem to be a non-negative (non-zero) KKT point if the objective value at that KKT point is non-negative (non-zero).
Problem setup: We adopt a supervised learning framework for training, where we assume $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$ is the training dataset, and let $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]\in\mathbb{R}^{d\times n}$ and $\mathbf{y}=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n}$. For a neural network $\mathcal{H}$, its output is denoted by $\mathcal{H}(\mathbf{x};\mathbf{w})$, where $\mathbf{x}$ is the input and $\mathbf{w}\in\mathbb{R}^{k}$ is the vector containing all the weights, and $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$ denotes the gradient of $\mathcal{H}(\mathbf{x};\mathbf{w})$ with respect to $\mathbf{w}$. We let $\mathcal{H}(\mathbf{X};\mathbf{w})=\left[\mathcal{H}(\mathbf{x}_{1};\mathbf{w}),\ldots,\mathcal{H}(\mathbf{x}_{n};\mathbf{w})\right]^{\top}\in\mathbb{R}^{n}$ be the vector containing the output of neural network for all inputs, and $\mathcal{J}(\mathbf{X};\mathbf{w}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ denotes the Jacobian of $\mathcal{H}(\mathbf{X};\mathbf{w})$ with respect to $\mathbf{w}$.
The following assumptions formalize the properties of neural networks considered in this paper.
@@
We make the following tripartite assumption on the neural network function: (i) For any fixed $\mathbf{x}$, $\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz and is definable under some o-minimal structure that includes polynomials and exponential. [^2] We note that definability in some o-minimal structure is a mild technical assumption and is satisfied by all modern deep neural networks. It allows us to ensure convergence of certain gradient flow trajectories (Ji & Telgarsky, 2020). (ii) For all $c>0$, $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),$ for some $L>2$ (i.e., the network is $L$-homogenous). (iii) The gradient $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz.
@@
where $\ell(\hat{y},y)$ is the loss function, and $\nabla_{\hat{y}}\ell(\hat{y},y)$ is assumed to be locally Lipschitz in $\hat{y}$, which is satisfied by typical loss functions such as square and logistic losses. For $\mathbf{z},\mathbf{y}\in\mathbb{R}^{n}$, we define $\ell^{\prime}(\mathbf{z},\mathbf{y})=[\nabla_{\hat{y}}\ell(z_{1},y_{1}),\ldots,\nabla_{\hat{y}}\ell(z_{n},y_{n})]^{\top}\in\mathbb{R}^{n}$. We minimize eq. 1 using gradient flow, which leads to the following differential equation
@@
where $\delta$ is a positive scalar that controls the scale of initialization, and $\mathbf{w}_{0}$ is a vector.
Following Kumar & Haupt (2024), for a fixed vector $\mathbf{z}\in\mathbb{R}^{n}$ and neural network $\mathcal{H}$, the Neural Correlation Function (NCF) is defined here as
@@
and measures the correlation between the vector $\mathbf{z}$ and the output of the neural network. The constrained NCF refers to the following constrained optimization problem
$\max_{\|\mathbf{w}\|^{2}_{2}=1}\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{w}).$
4 Early Directional Convergence
4.1 Main Result
The following theorem establishes the approximate directional convergence in the early stages of training $L$-homogeneous neural networks with small initializations.
Theorem 1.
Let $\mathbf{w}_{0}$ be a fixed unit vector. For any $\epsilon\in(0,\eta/2)$, where $\eta$ is a positive constant, there exists $T_{\epsilon},\tilde{B}_{\epsilon}$ and $\overline{\delta}>0$ such that the following holds: for any $\delta\in(0,\overline{\delta})$ and solution $\mathbf{w}(t)$ of eq. 2 with initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, we have
$\left\|\mathbf{w}\left({t}\right)\right\|_{2}\leq\tilde{B}_{\epsilon}\delta,\text{ for all }t\in\left[0,{T_{\epsilon}}/{\delta^{L-2}}\right].$